### v1

Abstract
We present a unified dataset for document Question-Answering (QA), which is obtained combining several public datasets related to Document AI and visually rich document understanding (VRDU). Our main contribution is twofold: on the one hand we reformulate existing Document AI tasks, such as Information Extraction (IE), into a Question-Answering task, making it a suitable resource for training and evaluating Large Language Models; on the other hand, we release the OCR of all the documents and include the exact position of the answer to be found in the document image as a bounding box. Using this dataset, we explore the impact of different prompting techniques (that might include bounding box information) on the performance of open-weight models, identifying the most effective approaches for document comprehension.
keywords:
Large Language Models (LLMs), Document AI, Dataset, Question Answering, Fine-tuning, Information Extraction.
1 Introduction
The increasing number of documents produced in various fields, including scientific research, legal proceedings, healthcare, and business, has created an enormous demand for efficient information extraction (IE) methods.
In document processing research, Optical Character Recognition (OCR) has proven essential for transforming scanned documents and images into machine-readable text, facilitating further analysis. Initially, statistical methods [1] were used alongside OCR to extract information, followed by machine learning approaches. Subsequently, deep learning techniques [2], especially methods related to natural language processing (NLP), became crucial in advancing document understanding. Today, the focus has shifted towards Large Language Models (LLMs) [3], which, with their exceptional ability to model natural language in complex contexts, have further enhanced document comprehension and the automation of information extraction from extensive volumes of text.
OCR tools and LLMs are now extensively used to perform several tasks in Document AI, including:
  • Document Image Classification: classifies document images into types such as invoices, scientific papers, and receipts [4].
  • Layout Analysis: examines a document’s structure, identifying elements like text, images, and tables [5];
  • Visual Information Extraction: extracts entities and relationships from unstructured content, considering text, visual elements, and layout [6];
  • Visual Question Answering: answers natural language questions based on a document’s content [7];
The two main motivations for building the BoundingDocs dataset [^1] The dataset is publicly available at https://huggingface.co/datasets/letxbe/BoundingDocs., that is focused on Information Extraction and Question Ansering, are:
  1. 1.
    the lack of extensive and diverse QA datasets in the field of Document AI;
  2. 2.
    the lack of precise spatial coordinates in the existing datasets.
Current datasets do not effectively incorporate positional data, which is essential for reducing hallucinations and improving performance by enabling LLMs to understand document layout more precisely.
Contribution
In this work, we propose a unified approach to build a Question-Answering dataset. Such a dataset can be used for evaluating how good Document AI models are to extract relevant information when answering to natural language questions. In doing so, we aim to address the following research questions:
  • RQ1: How can existing datasets be unified into a common Question-Answering format?
  • RQ2: Can rephrased questions generated by LLMs enhance answer accuracy for document-based questions?
  • RQ3: Does including layout information in prompts (e.g. [8, 9]) improve the model’s performance on document comprehension tasks?
To explore these questions, our study is organized into the following sections. Section 2 reviews the existing literature and benchmarks in the field of Document AI and question answering tasks; Section 3 describes the process of unifying datasets into a common Question-Answering format with enhanced layout annotations; Section 4 evaluates the performance of LLMs using various prompting techniques and presents the results. Conclusions are drawn in Section 5 where we discuss our findings, the key challenges encountered, and propose directions for future research.
2 State of the art
We provide an overview of the main models and techniques proposed for QA and Visual Question-Answering (VQA) [7]. We also discuss the features of the main datasets in the Document AI that we considered in our research.
2.1 Related Datasets
As summarized in Table 1, we selected datasets that best match our focus on comprehensive document understanding and advanced VQA, addressing challenges across both single-page and multi-page documents. For a more detailed review of datasets specific to Document Layout Analysis, please refer to our additional survey [10], which includes more datasets focused on layout-related tasks.
Dataset review
Dataset Size Answers OCR Info OCR Engine Type Lang
VRDU [11] 2,556 Yes 1 0 1 1
Deepform [12] 60,000 Yes 1 1 1 1
DUDE [13] 4,974 Yes 1 1, 2, 4 1 1
FATURA [14] 10,000 Yes 2 5 2 1
SP-DocVQA [7] 12,767 No 1 3 1 1
MP-DocVQA [15] 5,929 No 1 2 1 1
FUNSD [16] 199 Yes 1 0 1 1
Kleister Charity [17] 2,788 No 3 1, 2 1 1
Kleister NDA [17] 540 No 3 1, 2 1 1
SROIE [18] 1,000 No 1 0 1 1
XFUND [19] 1,393 Yes 1 0 1 2,3,4,5,6,7,8
SynthTabNet [20] 600,000 Yes 1 5 2 1
CORD [21] 1,000 Yes 2 0 1 9
GHEGA [22] 246 Yes 1 0 1 10
Table 1: Datasets review details. For clarity, the following codes are used in the table: OCR info - 1: Full text with bboxes, 2: Partial text with bboxes, 3: Full text without bboxes; OCR engine - 0: Not specified, 1: Tesseract, 2: Amazon Textract, 3: Microsoft OCR, 4: Azure Cognitive Service, 5: Synthetic document (OCR not needed; text is pre-known); Type - 1: Real, 2: Synthetic; Lang - 1: English, 2: Italian, 3: French, 4: Spanish, 5: Chinese, 6: German, 7: Portuguese, 8: Japanese, 9: Indonesian, 10: Not specified mix.
Among the foundational datasets, DocVQA [7, 15] stands as one of the earliest benchmarks dedicated to VQA on document images, focusing on understanding both textual and layout aspects of documents. Launched in 2020, DocVQA comprises multiple tasks designed to push the boundaries of document comprehension. The primary tasks include answering questions about individual document pages and analyzing multi-page documents — a crucial capability for real-world applications. The Single Page [7] subset of this dataset includes 50,000 questions over 12,767 documents, while the Multi Page [15] subset contains 46,436 questions spanning 5,929 documents (covering 47,952 pages in total). These datasets require models to interpret the visual structure of documents and to derive insights that go beyond simple text extraction.
DUDE [13] builds on this foundational work by extending VQA to multi-domain, multi-purpose documents. The dataset provides 5,000 annotated PDF files with 18,700 question-answer pairs across various domains and time frames, making it a unique resource for tasks that merge Document Layout Analysis with complex, layout-based question answering. Unlike typical QA datasets, DUDE often requires multi-step reasoning, handling both content and structural queries. For instance, questions may include layout-based prompts such as “How many text columns are there?” or require arithmetic and comparison skills, presenting a challenging dataset for models trained primarily on text-based QA.
Another significant resource is Docmatix [23], developed by HuggingFace and released during our research period. Docmatix introduces a vast dataset with $2.4$ million images and $9.5$ million question-answer pairs from $1.3$ million PDF documents, making it one of the largest publicly available DocAI resources. Generated from the PDFA dataset, Docmatix uses OCR-extracted text to produce diverse QA pairs via an automated approach, offering comprehensive coverage of document types and layouts. This dataset provides only document images with paired QA responses, excluding the original OCR text, which shifts focus toward layout and image-based comprehension.
In addition to these datasets, several others serve as standard benchmarks and are worth mentioning briefly. VRDU [11] includes two corpora—registration forms from the U.S. Department of Justice and ad-buy forms from the FCC—representing templates of varying complexity. The FATURA [14] dataset provides 10,000 images across 50 templates with imbalanced distributions for fields commonly found in invoices, such as buyer information and total amount, along with bounding box annotations for structured data extraction. Kleister [17] datasets offer specialized financial reports and legal documents, with Kleister Charity and Kleister NDA addressing entity extraction for key attributes. Deepform [12] offers approximately 20,000 labeled receipts for political ad purchases with labeled fields for specific political advertising details.
Finally, FUNSD [16] and XFUND [19] are form-centric datasets focused on entity linking and key-value extraction in noisy, often multilingual documents. FUNSD includes 199 annotated forms in English, designed for form understanding, while XFUND broadens this to a multilingual setting with documents in seven languages, capturing the diversity of form structures globally.
2.2 Related methods
In recent years, the QA task [7, 15] has been approached in many ways, leveraging different techniques and model architectures. These methods can be broadly categorized into NLP-based, LLMs, and multimodal architectures, each addressing different aspects of document understanding and question answering.
NLP-based approaches build on general Question-Answering models, primarily focusing on text semantics without explicitly incorporating document layout or visual features. A prime example is BertQA [7], which utilizes a BERT architecture followed by a classification head to predict the start and end indices of an answer span. Modifications such as changes in hyperparameters and the introduction of new pre-training tasks have been explored in multiple works [24, 25], resulting in improved outcomes.
LLM-based methods leverage large language models to perform document understanding tasks by encoding structural and layout information directly into the input. For instance, LMDX [26] incorporates layout information via bounding box coordinates in the prompt, enhancing retrieval precision and reducing hallucination. DocLLM [27], which builds on the LayoutLM family, includes a specialized pretraining phase focused on structured layout data to improve document layout understanding. In contrast, NuExtract [28] is designed for extracting structured JSON data from documents, using training data derived from the Colossal Clean Crawled Corpus [29].
Multimodal architectures combine visual and textual features to enhance document comprehension across layout, content, and structure. Among OCR-free methods, mPLUG-DocOWL 1.5 [30] integrates a Vision Transformer (ViT) [31] with an LLM for comprehensive Document AI analysis, aligning layout and textual cues effectively without requiring separate OCR stages. Similarly, Donut [32] and Dessurt [33] operate without OCR preprocessing, directly integrating image and text data for robust document understanding.
In contrast, OCR-dependent models further refine document comprehension by incorporating OCR-based tokens. Hi-VT5 [15], for example, combines OCR tokens with visual features, optimizing its effectiveness for Question-Answering tasks that rely on precise textual information. Additionally, LayoutLMv3 [34] introduces visual patch embeddings in place of traditional CNNs to better align text, layout, and visual cues, resulting in improved performance on tasks requiring fine-grained structural interpretation.
3 Dataset construction
Figure 1: Dataset construction pipeline. The rectangles represent processes while the parallelograms represent foundation LLM models.
We base our new dataset, BoundingDocs, on the following datasets selected from Table 1: SP-DocVQA, MP-DocVQA, DUDE, Deepform, VRDU, FATURA, Kleister Charity, Kleister NDA, FUNSD, and XFUND. This collection encompasses a diverse range of document types, linguistic features, and question-answer formats, providing essential resources for training and evaluating advanced Document AI models.
In Figure 1 we show the implemented pipeline for dataset construction.
3.1 Dataset format definition
For each document, a JSON file contains the annotation (examples in Figure 2). Each word in the answer is linked to its corresponding bounding box. Following established practices in the literature (e.g., LayoutLM [34], BERT [2]), the bounding boxes are normalized integers ranging from 0 to 1000 relative to the actual page size. Each bounding box is defined by a list of four values: the width, the height, the $X$ and $Y$ coordinates of the top-left vertex of the rectangle.
3.2 Producing annotations
A significant challenge comes from integrating various types of annotations into a unified structure. Datasets like Deepform, Kleister, and FATURA provide annotations that only establish a relationship between a key and its corresponding value in the text, such as annotating Address = 48 Woodford, SandyFord. However, these datasets lack essential positional information, such as the text’s location, frequency of occurrence, and page number. In contrast, datasets like VRDU and DocVQA include positional details that specify where the key value appears in the text. However, inconsistencies may arise because these datasets utilize different OCR tools, leading to variations in positional measurements and formats. To ensure consistent calculations for bounding box positions, Amazon Textract [35]. has been selected for this purpose.
In the case of FUNSD and XFUND, the datasets contain annotations related only to the text’s structure and relationships between elements. Consequently, additional steps are necessary to generate relevant questions from these datasets.
3.2.1 Dataset preparation
Upon collecting and downloading the datasets, the following preliminary operations have been considered case by case. These additional steps are critical to standardize and prepare the datasets for the generation of annotations.
Annotation Conversion: When the annotations in a dataset have an undocumented or complex format, they are converted into a standardized, more straightforward format. This is particularly necessary for the VRDU dataset, where the original annotations require interpretation and conversion.
Filtering Pages/Questions: Some datasets contain redundant or irrelevant content, as unnecessary pages or questions, that have been removed. For instance, in the DocVQA dataset, pages from the Multi Page set were excluded from the Single Page set to prevent duplication. Additionally, for both DUDE and DocVQA datasets, pure visual questions, i.e., lacking the answer as recognized by the OCR in the image, are filtered out.
Downloading Original Documents: In datasets where only annotations are provided without the corresponding documents, the original documents are downloaded from external sources. This step was necessary for the Deepform dataset, where the PDFs were not included alongside the annotations.
OCR Processing with Textract: To ensure consistency across all datasets, Amazon Textract has been applied to all documents, regardless of whether they already contained OCR data. Datasets were processed through Textract not only when OCR data was completely absent, but also when OCR was only provided for the annotated fields. This process has been applied to datasets such as VRDU, FATURA, Kleister, SP-DocVQA, Deepform, FUNSD, and XFUND, where OCR data is either insufficient or not provided.
Key-Value Association Creation: For certain datasets, key-value pairs for information extraction were manually generated from the annotations. For instance, in FUNSD and XFUND datasets the key-value associations are automatically created from existing document annotations. This step involves linking elements labeled as questions to their corresponding answers to facilitate coherent information extraction.
3.2.2 Matching annotations and OCR
To match the answer to each question with the data extracted by Textract [35], a script has been developed: the main challenge is to identify the correct word when the same value appears at multiple positions. A considerable time has been devoted to produce high quality annotations. This script, a significant part of our contribution, is used across all datasets with only slight modifications to match the different annotation formats.
For a document and a given key-value pair, where the key represents a label (such as “name,” “address,” or “date”) describing the type of information, and the value contains the actual data associated with that label, the script executes the following steps:
Deepform QA pair "deepform/8385": {     "question": "What is the Gross Amount?",     "answers": [         {             "value": "\$576,405.00",             "location": [ [90, 11, 364, 768] ],             "page": 1         }     ],     "rephrased_question":         "What is the value of the Gross Amount?" }
Kleister Charity QA pair "kleister_charity/73938": {     "question": "What is the Address Postcode?",     "answers": [         {             "value": "ST4 8AW",             "location": [ [34, 10, 692, 335] ],             "page": 1         }     ],     "rephrased_question":         "What is the postal code of the address?" }
Figure 2: Sample of QA pairs from the dataset. The left QA pair is sourced from Deepform, while the right one is from Kleister Charity. The purple values represent the specific details related to each QA pair, and the blue keys denote the fixed structure defined for our dataset.
  1. 1.
    Compare each text line extracted by Textract (Line) with the correct answer using Jaccard similarity. The Jaccard similarity between two sets $A$ and $B$ is given by: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ where $|A\cap B|$ is the number of common elements between the two sets, and $|A\cup B|$ is the total number of unique elements across both sets.
  2. 2.
    If similarity exceeds a given threshold, the Line is added to a set of candidates.
  3. 3.
    For each candidate line, verify that each word within it is also detected as a Word block by Textract and falls within the Line bounding box. These words and their positions form the extracted answer for each key.
  4. 4.
    Questions are generated using the template What is the [key name]? (e.g., What is the Address?). For XFUND, this template was translated to match document languages. Datasets with pre-defined questions (DUDE, MP-DocVQA, SP-DocVQA) used their own questions.
  5. 5.
    Moreover, for VRDU Ad Buy Form, additional questions are created to account for key-value pairs linked to specific ad programs, such as:
    • What is the [program_start_date] for [program_desc]?
    • What is the [program_end_date] for [program_desc]?
    • What is the [sub_amount] for [program_desc]?
3.2.3 Rephrasing questions
After completing the matching between annotations and OCR, the questions for the new dataset are generated. Inspection of these questions, which followed a simple template-based structure, revealed that they are often grammatically incorrect, overly simplistic, and consistently adhered to the same pattern. This raised concerns that training an LLM on these questions could introduce bias, potentially leading to poor performance on questions written by humans, which may not follow the template.
To mitigate this issue, we employed the Mistral 7B model [36] to correct and rewrite the questions, aiming to fix errors and introduce linguistic diversity. Other Mistral models, such as Mistral Large [37] and Mixtral 8x7B [38], were also tested, but they produced overly complex, verbose, and unnatural questions.
The prompt for question rewriting included manually written examples to guide the model, with no information about the correct answer to avoid biasing the generation. For example, the question What is the Gross Amount? was rewritten by the LLM as What is the value of the Gross Amount?.
This procedure was applied to most questions in the dataset, adding a new attribute, rephrased_question. Questions from DUDE, MP-DocVQA, and SP-DocVQA were excluded as they were already human-written. Additionally, questions from XFUND were excluded due to concerns over the model’s ability to generate questions in languages other than English.
Dataset Documents Pages Questions Ques./Page Ques./Doc
Deepform 24,345 100,747 55,926 0.55 2.30
DUDE 2,583 13,832 4,512 0.33 1.75
FATURA 10,000 10,000 102,403 10.24 10.24
FUNSD 199 199 1,542 7.75 7.75
Kleister Charity 2,169 47,550 8,897 0.19 4.10
Kleister NDA 337 2,126 696 0.33 2.07
MP-DocVQA 5,203 57,643 31,597 0.55 6.07
SP-DocVQA 266 266 419 1.58 1.58
VRDU Ad Form 641 1,598 22,506 14.08 35.11
VRDU Reg. Form 1,015 2,083 3,865 1.86 3.81
XFUND 1,393 1,393 16,653 11.95 11.95
Total 48,151 237,437 249,016 1.05 5.17
Table 2: Overall dataset statistics.
In Figure 2 you can observe an example of the final format of the dataset questions, including the rephrased version of the questions.
3.3 Statistics & splits
The dataset is split into training, validation, and test sets, using an 80-10-10 split based on document count, where all questions related to a single document are contained within the same set. Table 2 gives an overview of the dataset’s size and sources distribution. Detailed statistics can be found in the Supplementary Material.
To ensure that question types and document layouts are uniformly distributed across the three sets, documents from each source dataset are sampled separately. Specifically, documents from Deepform are split in an 80-10-10 ratio, followed by documents from FATURA, DUDE, and all the others. The union of these individual splits yields the final training, validation, and test sets, with balanced layout and document types across all sets.
Some of the pages annotated using the proposed algorithm and belonging to BoundingDocs are shown in figures 3 and 4. For illustration purpose colored rectangle is drawn around the fields corresponding to the correct answers to the questions.
Figure 3: Deepform page with bbox annotations.
Figure 4: VRDU Registration Form page with bbox annotations.
3.4 Dataset examples
Template Question Rephrased Question Answer
What is the Advertiser? Who is the advertiser? OBAMA FOR AMERICA
What is the Gross Amount? What is the value of Gross Amount? \$119,000.00
What is the Registrant Name? What is the name of the registrant? Greenfield & Kress P.A.
What is the Registration Number? What is the registration number for the company? 6294
Table 3: QA pairs of the examples in Figure 3 and Figure 4. The first two refer to the Deepform sample and the last two to the VRDU Registration Form one.
Figure 5: Finetuning and testing pipeline. The rectangles represent processes while the parallelograms represent foundation LLM models.
In Fig. 3 (Deepform) and Fig. 4 (VRDU Registration Form) it is possible to observe two pages while Table 3 contains their QA pairs. In Fig. 3 the extracted fields are the advertiser’s name and the gross amount for the various transmissions. In Fig. 4 the fields to be extracted are only two: the registrant name and the registration number.
These two examples illustrate how, despite the high number of documents in the collection, the potential amount of information present in the documents is underutilized, as the annotated fields are few compared to the entire body of the documents, indicating that the potential of this large document collection is not being properly exploited.
Additional examples that provide a full overview of the entire variety of the dataset can be found in the Supplementary Material.
4 Experimental Results
Table 4 presents our experimental results across the different datasets and model configurations. The finetuning and testing pipeline implemented is summarized and plotted in Figure 5.
Model Deepform DUDE FATURA FUNSD XFUND SP-VQA
Mistral-7B-v0.3$\star$ 42.3 9.1 6.8 14.3 6.1 22.2
0.22% 16.19% 1.03% 1.23% 15.85% 10.00%
[…]
0.47% 5.52% 0.12% 3.68% 9.55% 2.50%
[…]
6.79% 64.76% 7.40% 5.52% 2.07% 52.50%
Template-Template 97.7 70.5 99.9 75.7 70.1 75.3
0.00% 0.00% 0.00% 0.61% 0.61% 0.00%
Template-Rephrased 96.8 70.9 91.5 71.1 68.8 70.2
3.75% 1.14% 0.23% 0.00% 1.53% 2.50%
Rephrased-Template 97.7 70.4 99.7 71.8 67.6 76.6
0.00% 0.00% 0.00% 0.00% 0.67% 0.00%
Rephrased-Rephrased 97.1 71.2 99.8 72.3 68.2 76.1
0.00% 0.00% 0.00% 0.00% 0.92% 0.00%
Reph.-Reph.-bbox 97.7 73.4 99.3 78.8 71.2 82.1
5.01% 4.95% 5.97% 17.79% 10.34% 5.00%
Reph.-Reph.-bbox 97.7 72.1 99.3 74.7 70.3 83.0
w/ regex 0.80% 0.38% 4.50% 4.29% 0.98% 0.00%
Model Kl. Charity Kl. NDA MP-VQA VRDU-Ad VRDU-Reg. W. Avg.
Mistral-7B-v0.3$\star$ 21.2 32.5 12.5 23.1 28.6 22.4
4.84% 10.45% 7.86% 0.22% 1.52% 3.32%
[…]
2.93% 6.72% 3.54% 0.65% 6.09% 1.77%
[…]
2.07% 0.00% 52.82% 26.80% 4.31% 20.37%
Template-Template 91.9 66.3 75.5 96.7 96.5 91.3
0.00% 0.00% 0.06% 0.00% 0.00% 0.04%
Template-Rephrased 92.5 63.7 73.6 87.1 96.0 87.8
0.40% 1.49% 1.03% 0.13% 0.00% 1.62%
Rephrased-Template 91.8 61.1 73.6 96.2 95.2 90.7
0.00% 0.00% 0.06% 0.00% 0.00% 0.04%
Rephrased-Rephrased 92.3 64.4 73.3 96.4 95.7 90.6
0.00% 0.00% 0.07% 0.00% 0.00% 0.04%
Reph.-Reph.-bbox 92.8 61.6 76.0 96.4 96.7 91.6
4.34% 0.75% 7.16% 1.47% 0.51% 5.64%
Reph.-Reph.-bbox 92.9 61.9 75.8 96.1 96.8 91.3
w/ regex 0.03% 0.00% 0.35% 0.26% 0.00% 1.53%
Table 4: ANLS* scores and JSON parsing error percentages across datasets for each model in our custom dataset. ANLS* scores measure accuracy in answering document questions, while the bottom value in each cell shows JSON parsing errors, indicating output consistency. The first three rows list instruct models ($\star$); all remaining rows are fine-tuned versions of Mistral-7B-v0.3. Model names follow the ‘[training question type]-[testing question type]‘ format (e.g., ‘Template-Rephrased‘ means trained on template questions, tested on rephrased ones). ”bbox” indicates layout information is included in the prompt, and ”w/ regex” denotes that values were extracted with regex if JSON parsing failed. The ”W. Avg” column provides a weighted average across datasets, with bold and underlined values marking the top two scores per dataset.
4.1 Evaluation Metrics
We use the standard metric ANLS* [39], which supports a wider range of tasks including line-item extraction and document-processing tasks.
For each model-dataset pair in our results, we report two key measurements: the ANLS* value (rescaled between 0 and 100 for easier reading) and the percentage of non-JSON parsable responses relative to the total number of queries. The weighted average provides a comprehensive overview based on the number of examples for each dataset. For ANLS*, higher values indicate better performance, while for non-parsable responses, lower percentages are preferable. In our results presentation, the best values for each dataset are bolded, and second-best values are underlined.
4.2 Prompt Construction
In this study, each question in the dataset may have answers located on multiple pages. The significant computational costs associated with multi-page document processing, as reported e.g. by Multi PageDocVQA [15], together with the context size limitation of smaller LLMs, make us opt for an atomic approach instead of encoding everything in a single prompt.
For questions requiring information from multiple pages, we generate independent prompts for each relevant page, appending the same question to each prompt. For instance, if a five-page document contains relevant information on pages 2 and 4, we generate two prompts—one containing page 2’s content and the other page 4’s—each coupled with the question.
Each prompt includes the document text, the question, and a specification for the answer format (JSON), facilitating structured data extraction.
4.3 Baseline Models
We evaluated three popular open-weight models as baselines: Mistral 7B Instruct v0.3 [36], Llama 3 8B Instruct [40], and Phi 3.5 3.8B Instruct [41]. These models were chosen for their established performance and recognition in the NLP community. Testing was conducted on our custom dataset to establish initial benchmarks.
4.4 Ablation Study: Question Formulation
For investigating the impact of question formulation, we selected the Mistral 7B v0.3 (base version) for fine-tuning. We evaluated two types of questions—template-based (simple, consistent format) and rephrased (more varied, user-friendly language). Each model was tested with both question types, resulting in four experimental conditions:
  • Template-Template: Model trained and tested with template-based questions.
  • Template-Rephrased: Model trained with template-based questions, tested with rephrased questions.
  • Rephrased-Template: Model trained with rephrased questions, tested with template-based questions.
  • Rephrased-Rephrased: Model trained and tested with rephrased questions.
4.5 Incorporating Bounding Box Information
To assess the impact of spatial information, we incorporated bounding box coordinates into the prompts, denoted as Reph-Reph-bbox in Table 4. Each Textract-extracted text element in the prompt was annotated with bounding box coordinates, enabling the model to reference spatial context. In this configuration, the model was specifically fine-tuned to produce more complex JSON outputs that include not only the answer but also a comprehensive list of all locations where the extracted value appears in the document. While this approach provided richer spatial awareness, the requirement to generate more structured outputs introduced additional complexity that led to increased parsing errors.
To address these parsing challenges, we implemented the Reph-Reph-bbox w/regex configuration, which introduced a regex-based post-processing step. When the model’s structured JSON output was not parsable due to format inconsistencies or generation errors, the regex extraction mechanism served as a fallback solution to retrieve the target value, effectively maintaining the benefits of spatial information while mitigating the impact of parsing failures.
4.6 Research Question answers
Our experimental findings provide clear answers to our research questions, defined in Section 1:
  • RQ1 - Dataset Unification: By standardizing data from various sources (e.g., receipts, invoices, forms) into a consistent Question-Answering format, models are exposed to a wide range of document layouts and content types, enhancing their training efficiency. This unification significantly streamlines the fine-tuning process, making it easier to handle diverse document sources. Moreover, fine-tuned models show significant improvements over instruct models, quantitatively confirming that exposure to varied document formats and layouts enhances the model’s ability to extract information.
  • RQ2 - Question Formulation Impact: The study revealed significant insights into how different question formulation strategies affect document comprehension and information extraction. The Template-Template configuration demonstrated superior performance by leveraging structured, consistent question patterns. However, the Rephrased-Rephrased configuration emerged as a particularly robust solution, maintaining high ANLS* scores (e.g., 99.8 on FATURA, 96.4 on VRDU-Ad) while achieving 0% parsing errors across most datasets. This configuration showed remarkable versatility in handling both template-based and natural language queries. Notably, the Template-Rephrased setup performed least effectively, highlighting the challenges in transitioning from template-trained models to complex question structures.
  • RQ3 - Layout Information: The incorporation of spatial information in prompts yielded measurable improvements in model performance. The Reph-Reph-bbox configuration achieved the highest weighted average ANLS* (91.6) across all datasets, demonstrating consistent improvements over configurations without spatial information. Notable gains were observed in complex document understanding tasks, with ANLS* scores increasing to 71.2 on XFUND and 83.0 on SP-VQA. While the initial implementation showed increased parsing errors, the addition of regex-based post-processing (Reph-Reph-bbox w/regex) successfully maintained high performance while reducing error rates to competitive levels (1.53% weighted average).
5 Conclusions
The paper addresses the growing need for evaluating LLMs in Document AI tasks by proposing a unified dataset designed for document Question Answering taking into account the position of answers’ text in the document. Baseline experiments using open-weight LLMs demonstrate the challenges of applying generic models to specialized Document AI tasks; the performance of instruct models reveal clear limitations in generating correct and well-structured answers.
The paper reveals that while off-the-shelf LLMs struggle with document-specific tasks, targeted fine-tuning can significantly improve their capabilities. Rephrasing questions using LLMs improves the models’ understanding and response accuracy across different question formulations, suggesting that LLMs benefit from exposure to diverse linguistic variations during training. Incorporating layout and positional information into the prompt led to improved accuracy across most datasets, but at the cost of a higher percentage of non-parsable responses, reflecting the increased complexity of generating JSON outputs that include bounding box information.
In future work we aim to explore various methods for incorporating bounding boxes into prompts to better capture the spatial structure of documents and evaluate open-weight multimodal LLMs specifically designed to handle both textual and visual information. The constructed dataset and the experimental results provide a solid foundation for future research in Document QA. Fine-tuning models with enriched prompts has shown promising improvements.
Declarations
Competing interests The authors declare no competing interests.
Appendix A Dataset statistics
We now provide a quantitative illustration using tables and graphs to show the nature of the dataset in all its aspects.
Figure 6 and Figure 7 provide an overview of the dataset construction, showing how many documents and related questions from the various source datasets contribute to the overall dataset.
There are already several aspects to consider: first of all, it can be seen that Deepform is the dataset that contributes the most documents, but it has an average of about 2 questions per document, whereas the dataset that contributes the most questions is FATURA, with an average of more than 10 questions per document. Note that VRDU Ad Buy Form is the dataset that contains the most annotated fields, and both this aspect and the construction of additional questions for this particular dataset lead to a very high number of questions compared to the relatively low number of documents (an average of more than 34 questions per document). Also, note that there are very few documents related to SP-DocVQA: this is because, as already mentioned, most of the documents in this dataset were already present in MP-DocVQA, and there was no point in including them twice.
Figure 6: Documents and questions per dataset
Figure 7: Documents distribution across datasets
In Figure 8 and Table 5 you can observe the distribution of the languages in which the questions in the dataset are posed. The only questions, along with their respective documents, that are not in English are those formulated on XFUND, and thus they represent a clear minority compared to the total count.
Language Questions Percentage (%)
English 232,362 93.31
Italian 3,857 1.55
Spanish 2,753 1.11
French 2,176 0.87
German 2,564 1.03
Portuguese 3,743 1.50
Chinese 1,116 0.45
Japanese 445 0.18
Total 249,016 100.00
Table 5: Language distribution
Question Type Count Percentage (%)
Gross Amount 15848 28.34
Contract Number 7950 14.22
Flight From 7919 14.16
Flight To 7921 14.16
Advertiser 16288 29.12
Total 55926 100.00
Table 6: Question distribution for Deepform dataset
Question Type Count Percentage (%)
Buyer information 5653 5.52
Date of purchase 9800 9.57
Invoice ID 8796 8.59
Remarks and footers 5157 5.04
Seller Address 8131 7.94
Title 7346 7.17
Total amount after tax and discount 7992 7.80
Total words 3932 3.84
GSTIN 4708 4.60
To whom the invoice is sent 1356 1.32
Payment terms and conditions 2306 2.25
Discount 2383 2.33
Due date 5797 5.66
Seller email 4396 4.29
Total amount before tax and discount 6753 6.59
Tax 3799 3.71
Purchase order number 1400 1.36
Total amount to be paid 685 0.67
To whom the bill is sent 1285 1.25
Seller name 6728 6.57
Bank information 2600 2.55
Website of the seller 1400 1.38
Total 102403 100.00
Table 7: Question distribution for FATURA dataset
Question Type Count Percentage (%)
Charity Name 1617 18.17
Charity Number 2089 23.48
Spending Annually in British Pounds 66 0.74
Address Post Town 1948 21.90
Address Postcode 1621 18.22
Address Street Line 1495 16.80
Income Annually in British Pounds 61 0.69
Total 8897 100.00
Table 8: Question distribution for Kleister Charity dataset
Question Type Count Percentage (%)
Jurisdiction 319 45.83
Party 314 45.11
Term 62 8.91
Effective Date 1 0.15
Total 696 100.00
Table 9: Question distribution for Kleister NDA dataset
Question Type Count Percentage (%)
Gross Amount 614 2.72
Contract Number 615 2.73
Flight From 439 1.95
Flight To 440 1.96
Advertiser 584 2.59
Property 572 2.54
Agency 263 1.17
Product 561 2.49
Sub Amount 4197 18.65
Program Start Date 4714 20.95
TV Address 463 2.06
Channel 4556 20.24
Program End Date 4482 19.92
Program Description 6 0.03
Total 22506 100.00
Table 10: Question Distribution for VRDU Ad Buy Form dataset
Figure 8: Language distribution of questions
After providing a general overview of the dataset’s composition, it is also interesting to conduct an analysis of the types of questions that were generated and which field were extracted by running the matching algorithm on the various source datasets. Obviously, this analysis can only be conducted on the original datasets that pertain to key value extraction, as the questions are constructed according to the previously described template. For datasets such as DUDE and DocVQA, it is not possible to perform this type of tracking.
For Deepform, there are only five fields for which questions have been constructed, as can be observed in Table 6 and Figure 9. The main fields present are the total cost incurred for the advertisement and the name of the advertiser. The fields Flight From and Flight To are date values that represent the start and end days of the spot’s transmission.
Figure 9: Deepform question distribution
Regarding FATURA, the range of extracted fields is much broader compared to the previous Deepform, as visible in Figure 10 and Table 7. With 50 different layouts within FATURA, not all documents contain the same fields, which is the reason for the significant differences in frequencies among some fields. It is notable that the field with the most questions is Date of purchase (9800), while the least frequent is Total amount to be paid (685).
Figure 10: FATURA question distribution
Regarding the Kleister Charity dataset, the range of extracted fields is relatively narrow compared to the FATURA dataset. As shown in Table 8 and Figure 11, the dataset primarily focuses on extracting information such as the Charity Name, Charity Number, Address Post Town, Address Postcode, and Address Street Line. The fields with the fewest questions are Spending Annually in British Pounds and Income Annually in British Pounds, indicating that financial details are less frequently extracted from this dataset.
Figure 11: Kleister Charity question distribution
The Kleister NDA dataset, as detailed in Table and Figure 12, contains questions across a very limited set of fields: Jurisdiction, Party, Term, and Effective Date. The field with the most questions is Jurisdiction, followed by Party, while the Effective Date field has only a single question.
Figure 12: Kleister NDA question distribution
The VRDU Ad Buy Form dataset, as shown in Table 10 and Figure 13, contains a broader range of fields compared to the previous datasets. The fields with the most questions are Program Start Date, Channel, and Program End Date. In contrast, the field with the fewest questions is Program Description, with only 6 questions.
Figure 13: VRDU Ad Buy Form question distribution
The VRDU Registration Form dataset, as detailed in Table 11 and Figure 14, contains questions across 6 different fields. The fields with the most questions are Registration Number and Registrant Name, while the field with the fewest questions is Signer Title.
Question Type Count Percentage(%)
Registrant Name 959 24.81
Registration Number 983 25.43
File Date 783 20.25
Signer Name 654 16.93
Foreign Principle Name 264 6.83
Signer Title 222 5.75
Total 3865 100.00
Table 11: Question distribution for VRDU Registration Form dataset
Figure 14: VRDU Registration Form question distribution
The latest statistics worth noting are those related to the split made for training, validation, and testing. As previously described, the documents were divided according to an 80-10-10 percentage, assuming that the distribution of questions would be similar and that we would therefore obtain the same percentage division for the latter as well. As can be seen from the Table 12, our intuition was confirmed, achieving the desired partitioning for the questions as well.
Split Documents Questions % Documents % Questions
Train 38516 198601 80.0% 79.8%
Val 4804 24956 10.0% 10.0%
Test 4832 25463 10.0% 10.2%
Table 12: Train/Val/Test split
Appendix B Dataset examples
Similarly to what was done in the paper, a comprehensive qualitative overview of the entire variety of the dataset will be provided. An example will be shown for each source dataset, along with (almost) all the corresponding QA pairs formulated for that page, as shown in Table LABEL:tab:app_samples.
Table 13: QA pairs of the examples, each pair referencing the specific example it corresponds to.
Template Question Rephrased Question Answer Fig.
What is Advertiser? Who is the advertiser? Jordan, Jonathan 15
What is Gross Amount? What is the value of the Gross Amount? \$10,500.00 15
What is Address Post Town? What is the post town of the address? Stoke-on-Trent 16
What is Address Postcode? What is the postal code of the address? ST4 8AW 16
What is Address Street Line? What is the value of the Address Street Line? 28 Greenway 16
What is Charity Name? What is the name of the charity? Lucas’ Legacy - Childhood Brain Tumour Research 16
What is Charity Number? What is the charity number? 1167650 16
What is Jurisdiction? In which state is the company registered? Delaware 17
What is Party? What is the name of the company? Cisco Systems, Inc., 17
What is Contract ID? What is the contract ID number? 711207 18
What is the Product? What is the name of the product being advertised? Q42020 Broadcast 18
What is Property? What is the property name? KXLF 18
What is Agency? Who is the advertising agency? Left Hook Communications 18
What is Advertiser? Who is the advertiser? Bennett/Democrat/
Secretary of State
18
What is Gross Amount? What is the value of the gross amount? \$3,020.00 18
What is Sub Amount for M-F 530-7am News M-F 530-7am News? What is the value for the ’Sub Amount’ key for ’M-F 530-7am News’? \$100.00 18
What is the Channel for M-F 530-7am News M-F 530-7am News? What is the value of the ’Channel’ for the ’530-7am News’ broadcasted from Monday to Friday? All 18
What is Program Start Date for M-F 530-7am News M-F 530-7am News? What is the start date for the M-F 530-7am News program? 10/06/20 18
What is the Program End Date for M-F 530-7am News M-F 530-7am News? What is the end date for the program ’M-F 530-7am News’? 10/12/20 18
What is Registrant Name? What is the name of the registrant? KOREA TRADE PROMOTION CENTER 19
What is Registration Number? What is the registration number for the company? 1619 19
What is Signer Title? What is the signer’s title? DEPUTY DIRECTOR 19
First bubble in the HPA Axis? First bubble in the HPA Axis? Hypothalamus 20
What does CORT stand for in this document? What does CORT stand for in this document? Cortisol 20
Where does cortisol go after it is sent from the adrenal cortex? Where does cortisol go after it is sent from the adrenal cortex? Hypothalamus 20
What is Buyer information? What is the name of the buyer? Buyer :Nichole Harrington 8282 Kristie Lights South Loriburgh, PR 35228 US Tel:+(227)782-8066 Email:blackjames@
example.net Site:http://ruiz-bailey.com/
21
What is Date of purchase? When was the purchase date? Invoice Date: 30-Oct-1998 21
What is Due date? What is the due date? Due Date : 24-May-2020 21
What is Purchase order number? What is the purchase order number value? PO Number :72 21
What is Seller Address? What is the seller’s address? Address:05866 Velazquez Mount North Diane, NJ 20651 US 21
What is Total amount before tax and discount? What is the value of the total amount before tax and discount? SUB_TOTAL : 293.47 \$ 21
What is Tax? What is the tax amount? TAX:VAT (5.69%): 16.70 \$ 21
What is Title? What is the key for the title information? TAX INVOICE 21
What is Total amount to be paid? What is the value of the total amount to be paid? BALANCE_DUE : 305.39 \$ 21
What is MANUFACTURER:? What is the value of the manufacturer? R. J. REYNOLDS 22
What is BRAND NAME:? What is the value of the brand name? CARDINAL CIGARETTES (11 PACKINGS) 22
What is OTHER INFORMATION:? What is the value of OTHER INFORMATION? SEE ATTACHMENT 22
to whom is this letter written to? to whom is this letter written to? Mr. Rionda 23
when is the letter dated ? when is the letter dated ? October 18, 1940, 23
what is the auth. no. mentioned in the given form ? what is the auth. no. mentioned in the given form ? 5754 24
what is the value of percent per account as mentioned in the given form ? what is the value of percent per account as mentioned in the given form ? 50.06 24
what is the emp. no. mentioned in the given form ? what is the emp. no. mentioned in the given form ? 483378 24
what is the employee name mentioned in the given form ? what is the employee name mentioned in the given form ? IRENE KARL 24
what is the value of amount authorized per account ? what is the value of amount authorized per account ? 292.00 24
Qual è Cognome? Qual è Cognome? ANNI 25
Qual è Nome? Qual è Nome? GIACCOMO 25
Qual è Data Nascita? Qual è Data Nascita? 12/02/1988 25
Qual è Data? Qual è Data? 19/12/2020 25
Qual è Ora? Qual è Ora? 14:00 25
Figure 15: Deepform sample
Figure 16: Kleister Charity sample
Figure 17: Kleister NDA sample
Figure 18: VRDU Ad Buy Form. Not all the questions for this page are listed in Table LABEL:tab:app_samples, only until the details of the first broadcasting.
Figure 19: VRDU Registration Form sample.
Figure 20: DUDE sample.
Figure 21: FATURA sample.
Figure 22: FUNSD sample.
Figure 23: MP-DocVQA sample.
Figure 24: SP-DocVQA sample.
Figure 25: XFUND sample.